Suppose that X1 and X2 have the joint probability distribution

f(x1, x2) = kx1x2 x1 = 1, 2, 3 x2 = 2, 3, 4

0, otherwise

Find the value of k so that the two variables are independent of each other


1
Expert's answer
2022-05-10T23:15:28-0400

The joint probability density has to satisfy: "\sum_{x_1,x_2}f(x_1,x_2)=1". We will check that:

"\sum_{x_1,x_2}f(x_1,x_2)=\sum_{x_1=1,2,3;x_2=2,3,4}kx_1x_2=k(2+3+4+4+6+8+6+9+12)=54k". We receive that "k=\frac{1}{54}". Find the marginal densities: "f_{X_1}(x_1)=\sum_{x_2}f(x_1,x_2)=k(2x_1+3x_1+4x_1)=9kx_1".

"f_{X_2}(x_2)=\sum_{x_1}f(x_1,x_2)=k(x_2+2x_2+3x_2)=6kx_2".

As we can see: "f_{X_1}(x_1)f_{X_2}(x_2)=54k^2x_1x_2=kx_1x_2=f(x_1,x_2)". The latter implies that "X_1" and "X_2" are independent.